OpenStudy (.whitedragon.):

A survey found that 15 out of 100 people have blue eyes. On the basis of this survey, how many people in a group of 5000 are likely to have blue eyes?

OpenStudy (yanasidlinskiy):

15/100 = x /5000 [Set up as a ratio, sorry can't figure out the equation device on here!] .15 = x/5000 .15(5000) = x 750 = x

OpenStudy (anonymous):

Just use the relation 15/100 = x/5000 where x is the no. of people having blue eyesin a group of 5000 people x = 5000 x 15/100 = 750